# Relaxation and statistical equilibria in generalised two-dimensional flows

### v1

@@ 1 Introduction
The emergence of coherent structures in two-dimensional (2D) turbulence has long motivated the use of statistical mechanics as a theoretical framework, beginning with Onsager’s pioneering theory of point vortices (onsager1949statistical). Persistent vortices such as Jupiter’s Great Red Spot, the ubiquity of inverse energy cascades in the atmosphere and ocean, and the organisation of planetary flows into zonal jets, all highlight the tendency of geophysical systems to self-organise (byrne2013height; king2015upscale; shao2023physical; balwada2022direct; ScottWang2005; young2017forward). The large-scale dynamics of such systems are often modelled using the incompressible 2D Euler equations, which provide a natural setting for equilibrium statistical mechanics (miller1992statistical; bouchet2012statistical). In this framework, the large-scale coherent flow is interpreted as an emergent equilibrium state constrained by the conserved quantities of the system. Several studies have extended this approach to geophysically relevant flows (salmon1998lectures; majda2006nonlinear; bourouiba2008model; TeitelbaumMininni2012; weichman2022statistical).
@@
For the truncated two-dimensional (2D) Euler equations, kraichnan1975statistical derived the long time statistical equilibrium solutions using the canonical and microcanonical ensemble theory. For the truncated three-dimensional (3D) Euler equations, kraichnan1973helical showed that there are no negative temperature states. During relaxation toward equilibrium, subsets of Fourier modes thermalise and act as effective microstates, generating an effective viscosity effect on the modes that have not yet thermalised (cichowlas2005effective; bos2006dynamics). Moreover, studies have shown the truncated 3D Euler equations to reproduce aspects of Navier-Stokes dynamics at large scales (krstulovic2009cascades), and conversely, the large-scale flow of 3D Navier-Stokes equations also exhibit features characteristic of the truncated 3D Euler dynamics (Dallasetal2015; alexakis2019thermal).
@@
Understanding how a system relaxes toward equilibrium is a long-standing challenge in statistical mechanics. A central factor that controls relaxation toward equilibrium in such systems is the strength of the nonlinear interactions. The classical Fermi-Pasta-Ulam (FPU) problem exemplifies how weak nonlinearity and near-integrability can significantly delay relaxation, as energy is transferred only through weak wave-wave interactions (onorato2015route; onorato2023wave). In the weakly nonlinear regime, the timescale for equilibration can increase dramatically, resulting in slow or even incomplete thermalisation.
@@
The generalised 2D flows provide a systematic way to tune the effective nonlinearity of the governing equations through a parameter $\alpha$, which links the generalised vorticity $q$ and the streamfunction $\psi$ via $q=(-\nabla^{2})^{\alpha/2}\psi$ (pierrehumbert1994spectra). This formulation unifies a wide class of geophysical models: the barotropic vorticity equation for $\alpha=2$ (Tabeling2002; BoffettaEcke2012), the surface quasi-geostrophic (SQG) equation for $\alpha=1$ (held1995surface; Lapeyre2017), and a rescaled shallow-water quasi-geostrophic model for $\alpha=-2$ (LarichevMcWilliams1991; Smithetal2002). Recent work by jha2025cascades studied the transition in the direction of cascades of the inviscid conserved quantities as $\alpha$ varies in the forced-dissipative system. Positive values of $\alpha$ showed organisation of large scale vortices, while negative values of $\alpha$ showed the formation of filamentary structures reminiscent of 3D isotropic turbulence.
@@
In the limit $\alpha\to 0$, the nonlinear terms in the generalised model become progressively weaker and ultimately vanish, yielding a laminar flow state (jha2025cascades). This controlled reduction of nonlinearity naturally raises questions about how the dynamics and the relaxation to equilibrium of the inviscid system depend on $\alpha$, reminiscent in a way of the delayed thermalisation seen in the FPU model. Recent works on thermalisation in truncated 2D systems showed that for negative temperature states, where the energy is concentrated at small wavenumbers, the flows evolve to a partially thermalised state at long times (venaille2015violent; Agouaetal2025). Here, we focus on the long time thermalisation states of the generalised 2D model, and on understanding the dependence on the $\alpha$ parameter, as well as the limit of $\alpha\rightarrow 0$, where the nonlinearity tends to vanish.
@@ 2 Problem formulation and statistical equilibrium theory
Following pierrehumbert1994spectra, we consider the 2D advection equation for the generalised vorticity $q$ of an inviscid fluid,
@@
Equations (1) and (2) have two quadratic conserved quantities, namely,
@@
where $E_{G}$ is the generalised energy, $\Omega_{G}$ is the generalised enstrophy and the angular brackets $\langle{.}\rangle$ denote spatiotemporal averaging.
In Fourier space, relationships (2) and (3) become, respectively
@@
To derive the equilibrium states, we employ a canonical ensemble approach (lee1952; Kraichnan1967; salmon1998lectures), where the probability density of a given state is determined by the invariants (5) and (6) of the system and is given by
@@
where $\beta$ and $\gamma$ play the role of inverse temperatures, $\mathcal{Z}$ is the partition function, and $d\Gamma$ is the phase space volume element, which is given by $d\Gamma=\prod_{\bf k}d\Gamma(\mathbb{k})=\prod_{\bf k}d\widehat{\bf u}_{R}(\mathbb{k})\,d\widehat{\bf u}_{I}(\mathbb{k})$, such that $\widehat{\bf u}(\mathbb{k})=\widehat{\bf u}_{R}(\mathbb{k})+i\widehat{\bf u}_{I}(\mathbb{k})$.
Substituting definitions (5) and (6) into (9), we get
$\mathcal{Z}=\int\exp\left[-\beta\sum_{\bf k}k^{\alpha-2}|\widehat{\bf u}(\mathbb{k})|^{2}-\gamma\sum_{\bf k}k^{2\alpha-2}|\widehat{\bf u}(\mathbb{k})|^{2}\right]\prod_{\bf k}d\widehat{\bf u}_{R}(\mathbb{k})\,d\widehat{\bf u}_{I}(\mathbb{k})$ (9)
@@
The isotropic spectra of $E_{G}(k)$ and $\Omega_{G}(k)$ can be derived from the partition function using standard thermodynamic relations (kraichnan1975statistical; TeitelbaumMininni2012) as given below,
@@
These expressions reduce to well known equilibrium states for the 2D Euler ($\alpha=2$) and the SQG model ($\alpha=1$) as have been reported by kraichnan1975statistical and TeitelbaumMininni2012, respectively.
@@ 2.1 Equilibrium regimes and the $\alpha\mapsto-\alpha$ symmetry
Following Kraichnan1967; kraichnan1975statistical, the signs of $\beta$, $\gamma$ and the value of $k_{c}^{\alpha}=\Omega_{G}/E_{G}$ at $t=0$ determine the behaviour of the equilibrium spectra (14) distinguished by three regimes
@@
Regime III: $k_{\gamma}^{\alpha}<k_{c}^{\alpha}<k_{\max}^{\alpha},\quad$ $\beta>0,\quad-\beta<\gamma k_{\max}^{\alpha}<0.$ (17)
@@
Figure 1 demonstrates the three equilibrium regimes on the $\beta$-$\gamma$ parameter space for certain values of $\alpha$.
Boundaries of the regimes I and III are shown by dashed lines for the different values of $\alpha$.
@@
Regime I is characterised by $\beta<0$ and $\gamma>0$ (see Fig. 1). For $\alpha>0$, this implies that the generalised energy tends to accumulate toward $k_{\min}$, while generalised enstrophy is concentrated towards smaller scales (see Eq. (14)).
Regime II corresponds to $\beta>0$ and $\gamma>0$ (see Fig. 1), which is bounded by the energy-equipartition state ($\gamma=0$), $k_{c}=k_{\gamma}$, and the enstrophy-equipartition state ($\beta=0$), $k_{c}=k_{\beta}$, valid for all values of $\alpha$.
Finally, regime III is characterised by $\beta>0$ and $\gamma<0$ (see Fig. 1).
For $\alpha<0$, this implies that generalised enstrophy tends to accumulate toward $k_{\min}$, while generalised energy is concentrated towards smaller scales (see Eq. (14)). As kraichnan1975statistical noted, the negative temperature states exist only if $k_{\min}$ is non zero. Next, we proceed to show that the $\alpha\to-\alpha$ symmetry helps us understand the behaviour in more detail.
For the forced-dissipative evolution equation of the generalised vorticity $q$, the inverse cascade of $E_{G}$ and the forward cascade of $\Omega_{G}$ for $\alpha>0$, reverse direction when $\alpha<0$ (jha2025cascades). A natural question that arises is if such a behaviour is intrinsic for the ideal system (1), (2). Taking into account Eq. (2), the form of the invariants Eqs. (3) and the isotropic spectra Eqs. (14), we can deduce that if $\alpha\mapsto-\alpha$ then the solution is invariant under the interchanges $E_{G}(k)\leftrightarrow\Omega_{G}(k)$ and $\beta\leftrightarrow\gamma$. Since $\beta,\gamma$ are determined by the generalised energy and enstrophy, the symmetry is also given in terms of the interchange $E_{G}\leftrightarrow\Omega_{G}$. The exact expressions for $E_{G}$ and $\Omega_{G}$, are shown in Table 1 of Appendix A for certain values of $\alpha$, demonstrating the symmetry of $\alpha\mapsto-\alpha$. With this result, we can show that the solutions in regime I for $\alpha>0$ can be mapped to solutions in regime III for $\alpha<0$. Similarly the points in regime I for $\alpha<0$ can be mapped to points in regime III for $\alpha>0$, while solutions in regime II are mapped to other points in regime II.
@@
We fix the wavenumbers $k_{\min}=1$ and $k_{\max}=85$. The behaviour in regime I for different $\alpha$ is shown in Figure 2(a), where we fix $\beta=-10^{2}$ and $\gamma=10^{7}$. While for regime III, the behaviour for different $\alpha$ is shown in Fig. 2(b), where the values of $\beta=10^{7}$ and $\gamma=-10^{2}$ are exchanged compared to Figure 2(a). By comparing the figures, we can see the mapping from regime I to regime III as $\alpha\mapsto-\alpha$, the behaviour of the spectra $E_{G}(k)$ and $\Omega_{G}(k)$ are identical. The existence of this symmetry implies that it is sufficient to focus only on the positive $\alpha$ values.
Secondly, the exchange in behaviour between $E_{G}(k)$ and $\Omega_{G}(k)$ as $\alpha\mapsto-\alpha$ has similarities to the behaviour seen in the forced-dissipative system, where the cascade directions of the conserved quantities reverse when $\alpha\mapsto-\alpha$, (jha2025cascades).
@@ 2.2 The limit $\alpha\rightarrow 0$: multiple scale analysis
Since $J(\psi,\psi)=0$, this implies $\partial_{t}\psi=0$, i.e. the field is stationary and no energy transfer occurs. While at $\alpha=0$ the dynamics is trivial, in the limit of $\alpha\rightarrow 0$ we expect the dynamics to slow down due to the weak nonlinear term (jha2025cascades). To find the limiting behaviour of the system, we derive the leading order evolution equation that governs the dynamics of the generalised vorticity using the method of multiple scales.
We take $\alpha$ to be a small parameter, $0<\alpha\ll 1$, which sets the slow nonlinear time scale.
We then consider two different time scales, the fast timescale $t$ and the slow timescale $\tau=\alpha t$. Physically, the fast timescale can be thought as the eddy turnover time $t\sim 1/(Uk_{\min})$, where $U$ is the rms velocity. The field $\widehat{\psi}$ is then assumed to depend on both timescales as follows
@@ 3.1 Numerical set-up
To study the relaxation of the generalised vorticity equations towards equilibrium and the effect of $\alpha$ on the dynamics, we perform DNS in a periodic square domain of size $0,2\pi]\times[0,2\pi]$ of the truncated equations. We numerically integrate Eqs. (1), (2) using the pseudospectral method (GottliebOrszag1977) and a third-order Runge-Kutta scheme for the time advancement. The aliasing errors are removed with the 2/3 dealiasing rule, ensuring that the maximum wavenumber is $k_{\max}=N/3$ (mpicode2005), where $N$ denotes the number of aliased modes in the $x$ and $y$ directions. We fix the time-step to $dt=10^{-4}$ and we explore numerical resolutions $N^{2}=256^{2}$ and $512^{2}$. The flow is always initialised with a generalised energy spectrum $E_{G}(k)\propto k^{-2}$ for $k\in[1,4]$ and zero elsewhere. The most interesting statistical equilibrium solutions are the negative temperature states. For this reason, we focus our study to the solutions which are in regime I. With that in mind, at $t=0$ we fix the amplitude of the generalised energy $E_{G}(t=0)=1$ and $\Omega_{G}(t=0)$ is fixed by the energy spectrum of the initial conditions and the value of $\alpha$. For the values of $\alpha$ considered in this study, we find that $k_{c}=(\Omega_{G}/E_{G})^{1/\alpha}\in[1.0001,1.67]$. The values of $\beta$ and $\gamma$ for each $\alpha$ are obtained by numerically solving for $E_{G}(t=0)=\sum_{k}E_{G}(k)$ and $\Omega_{G}(t=0)=\sum_{k}\Omega_{G}(k)$. Tests were conducted for other initial conditions as well, and they are discussed in Appendix B.

### v2

@@ 1 Introduction
The emergence of coherent structures in two-dimensional (2D) turbulence has long motivated the use of statistical mechanics as a theoretical framework, beginning with Onsager’s pioneering theory of point vortices (Onsager, 1949). Persistent vortices such as Jupiter’s Great Red Spot, the ubiquity of inverse energy cascades in the atmosphere and ocean, and the organisation of planetary flows into zonal jets, all highlight the tendency of geophysical systems to self-organise (Byrne and Zhang, 2013; King et al., 2015; Shao et al., 2023; Balwada et al., 2022; Scott and Wang, 2005; Young and Read, 2017). The large-scale dynamics of such systems are often modelled using the incompressible 2D Euler equations, which provide a natural setting for equilibrium statistical mechanics (Miller et al., 1992; Bouchet and Venaille, 2012). In this framework, the large-scale coherent flow is interpreted as an emergent equilibrium state constrained by the conserved quantities of the system. Several studies have extended this approach to geophysically relevant flows (Salmon, 1998; Majda and Wang, 2006; Bourouiba, 2008; Teitelbaum and Mininni, 2012; Weichman and Marston, 2022).
@@
For the truncated two-dimensional (2D) Euler equations, Kraichnan (1975) derived the long time statistical equilibrium solutions using the canonical and microcanonical ensemble theory. For the truncated three-dimensional (3D) Euler equations, Kraichnan (1973) showed that there are no negative temperature states. During relaxation toward equilibrium, subsets of Fourier modes thermalise and act as effective microstates, generating an effective viscosity effect on the modes that have not yet thermalised (Cichowlas et al., 2005; Bos and Bertoglio, 2006). Moreover, studies have shown the truncated 3D Euler equations to reproduce aspects of Navier-Stokes dynamics at large scales (Krstulovic et al., 2009), and conversely, the large-scale flow of 3D Navier-Stokes equations also exhibit features characteristic of the truncated 3D Euler dynamics (Dallas et al., 2015; Alexakis and Brachet, 2019).
@@
Understanding how a system relaxes toward equilibrium is a long-standing challenge in statistical mechanics. A central factor that controls relaxation toward equilibrium in such systems is the strength of the nonlinear interactions. The classical Fermi-Pasta-Ulam (FPU) problem exemplifies how weak nonlinearity and near-integrability can significantly delay relaxation, as energy is transferred only through weak wave-wave interactions (Onorato et al., 2015, 2023). In the weakly nonlinear regime, the timescale for equilibration can increase dramatically, resulting in slow or even incomplete thermalisation.
@@
The generalised 2D flows provide a systematic way to tune the effective nonlinearity of the governing equations through a parameter $\alpha$, which links the generalised vorticity $q$ and the streamfunction $\psi$ via $q=(-\nabla^{2})^{\alpha/2}\psi$ (Pierrehumbert et al., 1994). This formulation unifies a wide class of geophysical models: the barotropic vorticity equation for $\alpha=2$ (Tabeling, 2002; Boffetta and Ecke, 2012), the surface quasi-geostrophic (SQG) equation for $\alpha=1$ (Held et al., 1995; Lapeyre, 2017), and a rescaled shallow-water quasi-geostrophic model for $\alpha=-2$ (Larichev and McWilliams, 1991; Smith et al., 2002). Recent work by Jha et al. (2025) studied the transition in the direction of cascades of the inviscid conserved quantities as $\alpha$ varies in the forced-dissipative system. Positive values of $\alpha$ showed organisation of large scale vortices, while negative values of $\alpha$ showed the formation of filamentary structures reminiscent of 3D isotropic turbulence.
@@
In the limit $\alpha\to 0$, the nonlinear terms in the generalised model become progressively weaker and ultimately vanish, yielding a laminar flow state (Jha et al., 2025). This controlled reduction of nonlinearity naturally raises questions about how the dynamics and the relaxation to equilibrium of the inviscid system depend on $\alpha$, reminiscent in a way of the delayed thermalisation seen in the FPU model. Recent works on thermalisation in truncated 2D systems showed that for negative temperature states, where the energy is concentrated at small wavenumbers, the flows evolve to a partially thermalised state at long times (Venaille et al., 2015; Agoua et al., 2025). Here, we focus on the long time thermalisation states of the generalised 2D model, and on understanding the dependence on the $\alpha$ parameter, as well as the limit of $\alpha\rightarrow 0$, where the nonlinearity tends to vanish.
@@ 2 Problem formulation and statistical equilibrium theory
Following Pierrehumbert et al. (1994), we consider the 2D advection equation for the generalised vorticity $q$ of an inviscid fluid,
@@
where $E_{G}$ is the generalised energy, $\Omega_{G}$ is the generalised enstrophy and the angular brackets $\langle{.}\rangle$ denote spatiotemporal averaging. In Fourier space, relationships (2) and (3) become, respectively
@@
To derive the equilibrium states, we employ a canonical ensemble approach (Lee, 1952; Kraichnan, 1967; Salmon, 1998), where the probability density of a given state is determined by the invariants (5) and (6) of the system and is given by
@@
where $\beta$ and $\gamma$ play the role of inverse temperatures, $\mathcal{Z}$ is the partition function, and $d\Gamma$ is the phase space volume element, which is given by $d\Gamma=\prod_{\bf k}d\Gamma(\mathbb{k})=\prod_{\bf k}d\widehat{\bf u}_{R}(\mathbb{k})\,d\widehat{\bf u}_{I}(\mathbb{k})$, such that $\widehat{\bf u}(\mathbb{k})=\widehat{\bf u}_{R}(\mathbb{k})+i\widehat{\bf u}_{I}(\mathbb{k})$. Substituting definitions (5) and (6) into (9), we get
$\mathcal{Z}=\int\exp\left[-\beta\sum_{\bf k}k^{\alpha-2}|\widehat{\bf u}(\mathbb{k})|^{2}-\gamma\sum_{\bf k}k^{2\alpha-2}|\widehat{\bf u}(\mathbb{k})|^{2}\right]\prod_{\bf k}d\widehat{\bf u}_{R}(\mathbb{k})\,d\widehat{\bf u}_{I}(\mathbb{k}).$ (9)
@@
The isotropic spectra of $E_{G}(k)$ and $\Omega_{G}(k)$ can be derived from the partition function using standard thermodynamic relations (Kraichnan, 1975; Teitelbaum and Mininni, 2012) as given below,
@@
These expressions reduce to well known equilibrium states for the 2D Euler ($\alpha=2$) and the SQG model ($\alpha=1$) as have been reported by Kraichnan (1975) and Teitelbaum and Mininni (2012), respectively.
@@ 2.1 Equilibrium regimes and the $\alpha\mapsto-\alpha$ symmetry
Following Kraichnan (1967, 1975), the signs of $\beta$, $\gamma$ and the value of $k_{c}^{\alpha}=\Omega_{G}/E_{G}$ at $t=0$ determine the behaviour of the equilibrium spectra (14) distinguished by three regimes
@@
Regime III: $k_{\gamma}^{\alpha}<k_{c}^{\alpha}<k_{\max}^{\alpha},\quad$ $\beta>0,\quad-\beta<\gamma k_{\max}^{\alpha}<0,$ (17)
@@
Figure 1 demonstrates the three equilibrium regimes on the $\beta$-$\gamma$ parameter space for certain values of $\alpha$. Boundaries of the regimes I and III are shown by dashed lines for the different values of $\alpha$.
@@
Regime I is characterised by $\beta<0$ and $\gamma>0$ (see Fig. 1). For $\alpha>0$, this implies that the generalised energy tends to accumulate toward $k_{\min}$, while generalised enstrophy is concentrated towards smaller scales (see Eq. (14)). Regime II corresponds to $\beta>0$ and $\gamma>0$ (see Fig. 1), which is bounded by the energy-equipartition state ($\gamma=0$), $k_{c}=k_{\gamma}$, and the enstrophy-equipartition state ($\beta=0$), $k_{c}=k_{\beta}$, valid for all values of $\alpha$. Finally, regime III is characterised by $\beta>0$ and $\gamma<0$ (see Fig. 1). For $\alpha<0$, this implies that generalised enstrophy tends to accumulate toward $k_{\min}$, while generalised energy is concentrated towards smaller scales (see Eq. (14)). As Kraichnan (1975) noted, the negative temperature states exist only if $k_{\min}$ is non zero. Next, we proceed to show that the $\alpha\to-\alpha$ symmetry helps us understand the behaviour in more detail.
For the forced-dissipative evolution equation of the generalised vorticity $q$, the inverse cascade of $E_{G}$ and the forward cascade of $\Omega_{G}$ for $\alpha>0$, reverse direction when $\alpha<0$ (Jha et al., 2025). A natural question that arises is if such a behaviour is intrinsic for the ideal system (1), (2). Taking into account Eq. (2), the form of the invariants Eqs. (3) and the isotropic spectra Eqs. (14), we can deduce that if $\alpha\mapsto-\alpha$ then the solution is invariant under the interchanges $E_{G}(k)\leftrightarrow\Omega_{G}(k)$ and $\beta\leftrightarrow\gamma$. Since $\beta,\gamma$ are determined by the generalised energy and enstrophy, the symmetry is also given in terms of the interchange $E_{G}\leftrightarrow\Omega_{G}$. The exact expressions for $E_{G}$ and $\Omega_{G}$, are shown in Table 1 of Appendix A for certain values of $\alpha$, demonstrating the symmetry of $\alpha\mapsto-\alpha$. With this result, we can show that the solutions in regime I for $\alpha>0$ can be mapped to solutions in regime III for $\alpha<0$. Similarly the points in regime I for $\alpha<0$ can be mapped to points in regime III for $\alpha>0$, while solutions in regime II are mapped to other points in regime II.
@@
We fix the wavenumbers $k_{\min}=1$ and $k_{\max}=85$. The behaviour in regime I for different $\alpha$ is shown in Figure 2(a), where we fix $\beta=-10^{2}$ and $\gamma=10^{7}$. While for regime III, the behaviour for different $\alpha$ is shown in Fig. 2(b), where the values of $\beta=10^{7}$ and $\gamma=-10^{2}$ are exchanged compared to Figure 2(a). By comparing the figures, we can see the mapping from regime I to regime III as $\alpha\mapsto-\alpha$, the behaviour of the spectra $E_{G}(k)$ and $\Omega_{G}(k)$ are identical. The existence of this symmetry implies that it is sufficient to focus only on the positive $\alpha$ values. Secondly, the exchange in behaviour between $E_{G}(k)$ and $\Omega_{G}(k)$ as $\alpha\mapsto-\alpha$ has similarities to the behaviour seen in the forced-dissipative system, where the cascade directions of the conserved quantities reverse when $\alpha\mapsto-\alpha$, (Jha et al., 2025).
@@ 2.2 The limit $\alpha\rightarrow 0$: multiple scale analysis
Since $J(\psi,\psi)=0$, this implies $\partial_{t}\psi=0$, i.e. the field is stationary and no energy transfer occurs. While at $\alpha=0$ the dynamics is trivial, in the limit of $\alpha\rightarrow 0$ we expect the dynamics to slow down due to the weak nonlinear term (Jha et al., 2025). To find the limiting behaviour of the system, we derive the leading order evolution equation that governs the dynamics of the generalised vorticity using the method of multiple scales.
We take $\alpha$ to be a small parameter, $0<\alpha\ll 1$, which sets the slow nonlinear time scale. We then consider two different time scales, the fast timescale $t$ and the slow timescale $\tau=\alpha t$. Physically, the fast timescale can be thought as the eddy turnover time $t\sim 1/(Uk_{\min})$, where $U$ is the rms velocity. The field $\widehat{\psi}$ is then assumed to depend on both timescales as follows
@@ 3.1 Numerical set-up
To study the relaxation of the generalised vorticity equations towards equilibrium and the effect of $\alpha$ on the dynamics, we perform DNS in a periodic square domain of size $0,2\pi]\times[0,2\pi]$ of the truncated equations. We numerically integrate Eqs. (1), (2) using the pseudospectral method (Gottlieb and Orszag, 1977) and a third-order Runge-Kutta scheme for the time advancement. The aliasing errors are removed with the 2/3 dealiasing rule, ensuring that the maximum wavenumber is $k_{\max}=N/3$ (Gómez et al., 2005), where $N$ denotes the number of aliased modes in the $x$ and $y$ directions. We fix the time-step to $dt=10^{-4}$ and we explore numerical resolutions $N^{2}=256^{2}$ and $512^{2}$. The flow is always initialised with a generalised energy spectrum $E_{G}(k)\propto k^{-2}$ for $k\in[1,4]$ and zero elsewhere. The most interesting statistical equilibrium solutions are the negative temperature states. For this reason, we focus our study to the solutions which are in regime I. With that in mind, at $t=0$ we fix the amplitude of the generalised energy $E_{G}(t=0)=1$ and $\Omega_{G}(t=0)$ is fixed by the energy spectrum of the initial conditions and the value of $\alpha$. For the values of $\alpha$ considered in this study, we find that $k_{c}=(\Omega_{G}/E_{G})^{1/\alpha}\in[1.0001,1.67]$. The values of $\beta$ and $\gamma$ for each $\alpha$ are obtained by numerically solving for $E_{G}(t=0)=\sum_{k}E_{G}(k)$ and $\Omega_{G}(t=0)=\sum_{k}\Omega_{G}(k)$. Tests were conducted for other initial conditions as well, and they are discussed in Appendix B.